OpenStudy (anonymous):

Solve the corresponding homogeneous system of equations using Gaussian elimination and back-substitution

OpenStudy (anonymous):

x-2y-3z=2 y-9z=5 -3x+7y=-1

OpenStudy (anonymous):

I solved the solution and got x=12 =0 and y=5 but is that what it's asking for?? I thought homogeneous meant that they equal 0

OpenStudy (anonymous):

z=0**

OpenStudy (phi):

you are correct. I think they want you to solve for the system = 0

OpenStudy (anonymous):

so I ignore what they are equal to?

OpenStudy (phi):

yes, you replace 2 5 -1 with 0 0 0

OpenStudy (anonymous):

okay perfect thanks that helps a lot!

OpenStudy (phi):

the solution to the homogenous system can be added to the solution you already found (because you are adding zero) the complete solution would be [ 12 5 0 ] + n* homogenous solution where n is an arbitrary constant

OpenStudy (anonymous):

In this example, wouldn't x,y and z equal zero?

OpenStudy (anonymous):

~_~

OpenStudy (anonymous):

I don't understand ok!!

OpenStudy (phi):

x-2y-3z=0 y-9z=0 -3x+7y=0 with just the coefficients 1 -2 -3 0 0 1 -9 0 -3 7 0 0 add 3 times the 1st row to the 3rd row 1 -2 -3 0 0 1 -9 0 0 1 -9 0 add -1 times the 2nd row to the 3rd row 1 -2 -3 0 0 1 -9 0 0 0 0 0 if we pick z=1, we can solve for x and y in terms of z if we continue to reduced row echelon form add 2 times 2nd row to 1st row 1 0 -21 0 0 1 -9 0 0 0 0 0 we can "read off" the solution vector as [21 9 1] the solution is z * [ 21 9 1 ] where z is any number this is what you should feet if you solved for x and y in terms of z up above